Rhesus factor, is a type of protein found on the outside of red blood cells. The protein is genetically inherited. A donor is said to have Rh+ blood if the Rhesus factor is present in their blood. Suppose that a blood donation campaign is launched by the University of Ottawa walk-in clinic. Among undergraduate students donating blood to walk-in clinic, 30% have Rh+. 1. If seven undergraduate students donate blood at this clinic on a particular day, what is the probability that at least three of them do not have the Rh factor? What is the probability that the number of students with Rh+ blood will be within 0.1 of the true population mean? 2. How many undergraduate students must donate blood to have the probability of at least one Rh+ donor over 0.90?
Answer #1

ho Let 2 p = 0.3 who do not have Rh factos. be the probability that an under graduate student, donating blood to the walk-in-Txo.3 , M = np. 2.1 We have P(14-HI<0-1) = P(-0.124-H 20.1) P(-0.1+2:1<4 <0.1+2.1) = P(2 <Y<2.2) such that, 2. We have to fin